OpenStudy (anonymous):

If sine of x equals 1 over 2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Written in more compact form, sin x = 1/2, or sin x = (1/2).

OpenStudy (mathmale):

You might want to sketch this situation, showing the lengths of all 3 sides of a right triangle with angle x such that sin x = 1/2. Use the DRAW feature, below.

OpenStudy (anonymous):

HMM ok so I got cos=-radical 3/2?

OpenStudy (anonymous):

I have the answers, I just want you to check them:) because I know u have to go soon

OpenStudy (mathmale):

Your cos=-radical 3/2 needs re-writing to be correct. You're discussing the cosine of an angle, but you don't show any angle (argument) here. Can you fix that?

OpenStudy (anonymous):

um +radical 3/2?

OpenStudy (mathmale):

No. Let's focus on the COSINE function. What is its argument? Go back to the original problem statement for a clue.

OpenStudy (anonymous):

um to find cos and tan?

OpenStudy (mathmale):

every trig function has an argument. we cannot write sin, cos, tan, etc., without arguments. What I'm trying to tell you is that your cos cannot stand alone. You mean cos x.

OpenStudy (mathmale):

Are you familiar with the word "argument" in this discussion?

OpenStudy (anonymous):

no:(

OpenStudy (mathmale):

you are working with trig functions. The "x" in cos x is your argument; another word for it would be "input" or "independent variable." \

OpenStudy (anonymous):

ok:)

OpenStudy (mathmale):

Your cos=-radical 3/2 must be re-written so that the cosine function has an argument. Please do that now.

OpenStudy (anonymous):

y= cos=-radical 3/2? We dont have to that for my teacher though

OpenStudy (mathmale):

It's VERY important that you do this. Again, give the cosine function an input OR argument OR independent variab le. What you want to write here is cos x = something.

OpenStudy (mathmale):

cos=-radical 3/2 becomes cos x = -radical 3/2. You should write the right side as -(radical 3)/2; if you don't, "radical 3/2" is ambiguous.

OpenStudy (mathmale):

So, you should end up with cos x =-(radical 3)/2

OpenStudy (mathmale):

Quiz: What does the "x" represent?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

then tan(x) would be radical 3/3 right?

OpenStudy (mathmale):

You're welcome. See you again!

OpenStudy (mathmale):

Again, J, you have to enclose "radical 3" in parentheses. Try again.

OpenStudy (anonymous):

no! Wait! Was I right about tan(x)??

OpenStudy (mathmale):

Also, please remember to insert the "argument" x next to the function "tan"

OpenStudy (anonymous):

(radical 3)/3?

OpenStudy (mathmale):

Your tan(x) was and is just fine. So, tan(x) = (radical 3)/3. Very good!

OpenStudy (anonymous):

Thanks! I have one last question though?

OpenStudy (mathmale):

Actually, J, tan (x) = (radical 3)/2. Go ahead with your last question.

OpenStudy (anonymous):

Using complete sentences, explain the key features of the graph of the cosine function

OpenStudy (anonymous):

Really? Are you sure? Before that, I got 1/(radical 3)

OpenStudy (mathmale):

I'm wrong. You're right. ;) tan(x) = 1/(radical 3), which can be rewritten as (radical 3)/3. Sorry.

OpenStudy (anonymous):

It's okay:)

OpenStudy (anonymous):

u didn't answer my last one though...?

OpenStudy (mathmale):

"Using complete sentences, explain the key features of the graph of the cosine function" I'd be glad to comment on YOUR remarks. What are characteristics of the graph of the cosine function?

OpenStudy (mathmale):

Can y ou draw the cosine function without reference to a book or to the Internet?

OpenStudy (anonymous):

um it has minimums and maximums?

OpenStudy (mathmale):

Yes, it does. So does the sine function. What else can you say about the graph of the cosine function? Draw it, if you haven't already.

OpenStudy (anonymous):

so cosine have a period of 2pi?

OpenStudy (mathmale):

Right. It also has repeated minima and repeated maxima, every 2Pi units along the x-axis. Anything else?

OpenStudy (mathmale):

Is the cosine function even or odd? How do you know?

OpenStudy (anonymous):

odd. And by guessing:o

OpenStudy (mathmale):

Better bone up on "even" and "odd." The cosine function is EVEN. Note that the graph is symmetrical with respect to the y axis.

OpenStudy (anonymous):

Ahh! i will do that:)

OpenStudy (mathmale):

R u now OK? Moe questions?

OpenStudy (anonymous):

Can we find 2 more facts about the cosine function?

OpenStudy (mathmale):

What is the max value? What is the min. value? What is the domain? What is the range? Because this function repeats itself (graph repeats itself), what is a classifier (name) for this and the sine and the tan functions?

OpenStudy (anonymous):

the domain could be all real #

OpenStudy (mathmale):

right, and the range?

OpenStudy (mathmale):

If you're not sure of the meaning of "range," please look it up (as it pertains to math)

OpenStudy (anonymous):

i know what it is, I jiust dont know what it could be

OpenStudy (mathmale):

Define "range."

OpenStudy (anonymous):

its like the output of the x values?

OpenStudy (anonymous):

so, 3x^2+12x+16 x=-2 y=-4

OpenStudy (mathmale):

The "range" of a function is an INTERVAL. What does INTERVAL mean in math, and what is the vertical INTERVAL pertaining to the cosine function?

OpenStudy (anonymous):

idek:/

OpenStudy (mathmale):

What is the minimum value of the cosine function? the max value?

OpenStudy (anonymous):

the maximum point it can be

OpenStudy (anonymous):

y=1?

OpenStudy (mathmale):

You mean "maximum value." Right. What is the minimum value the cos x function can have?

OpenStudy (anonymous):

0

OpenStudy (mathmale):

Actually, about half of the graph of the cosine function is below the x-axis; half is above it. Write the interval, using square brackets [ ], that represents the range of the cosine function.

OpenStudy (anonymous):

kk:)

OpenStudy (anonymous):

how would i find the minimum of g(x)=2*sin(x-pi)?

OpenStudy (mathmale):

Let's finish y = cos x first. What is the min. value? the max value? To answer your question: the min value is -2...but that's a separate question.

OpenStudy (anonymous):

alright! Sorry:(

OpenStudy (mathmale):

The min. value of the cos function is -1; the max is +1. Be certain that you know why this is and can explain it. Now use the numerals -1 and 1 to write the RANGE of the cos function.

OpenStudy (anonymous):

y=x>-1?

OpenStudy (mathmale):

That's an inequality. An interval looks like [a,b], (a,b], (a,b), etc. "a" represents the minimum, "b" the max. WRite the interval defined by the end points -1 and 1.

OpenStudy (anonymous):

i give up:(

OpenStudy (mathmale):

Min is -1; Max is +1. The range is the closed interval [-1,1].

OpenStudy (mathmale):

Please make a list of any vocabulary words with which you're not familiar. We could go over that after you've tried looking them up in a dictionary or in your math book or online.

OpenStudy (anonymous):

kk! I will do that:) Will u be on tomorrow at this time?

OpenStudy (mathmale):

Can't promise; probably won't be here because of a planned out of town trip. But you can message me and I can respond when back. I'd suggest you make a list of the questions you have and then try your best to find the answers before bringing up the questions with me. Best of luck to you. Bye for now!

OpenStudy (anonymous):

byee<3 thanks for all the help:)

OpenStudy (mathmale):

bye!